OpenStudy (anonymous):

How do I simplify: √2 + √5 / √6 - √10? The answer is supposed to be: - 5√2 + 2√3 + 2√5 + √30 / 4 but I get confused with the 3rd step. HELP!

OpenStudy (anonymous):

I know i have to multiply √2 + √5 / √6 - √10 by √6 + √10 / √6 + √10

OpenStudy (anonymous):

and then do I: √12 + √20 + √30 + √50 / √36 + √60 -√60 - √100?

OpenStudy (campbell_st):

\[2\sqrt{3}= \sqrt{12} \] \[\sqrt{20}=2\sqrt{5}\] \[\sqrt{50}=5\sqrt{2}\] since the donominator is -4 reverse the sign of each term in the numerator and and make the denominator positive

OpenStudy (anonymous):

Why is it: √20=√25?

OpenStudy (anonymous):

* √20 = 2√5?

OpenStudy (anonymous):

please answer...

OpenStudy (campbell_st):

\[\sqrt{20}=\sqrt{4\times5} =\sqrt{4}\times \sqrt{5}\]